OpenStudy (anonymous):

HELP!! Solve 2.42(4x − 6) = 2.12x − 2.45 for I have been trying to do this and im getting a wierd answer -.-

OpenStudy (anonymous):

for x

OpenStudy (precal):

2.42(4x-6)=1.12x-2.45 9.68x-14.52=1.12x-2.45

OpenStudy (precal):

8.56x=12.07 x=1.410046729

OpenStudy (anonymous):

it was wrong -.-

OpenStudy (precal):

no I checked it, it is right

OpenStudy (precal):

Were you told how to round?

OpenStudy (anonymous):

no but what is the answer round . maybe its right

OpenStudy (precal):

you have to be told how to round round to the first place 1.4 round to the second place 1.41

OpenStudy (anonymous):

its still wrong . what ever.

OpenStudy (precal):

Well, you might want to double check your instructions. This is the problem with inputing solutions online. The program will only say it is a valid solution if it is recorded correctly. the fractionsolution is \[\frac{1207}{856}\]